# Channel Coding for Unequal Error Protection in Digital Semantic Communication

### v1

@@ IV-B Bit Grouping Algorithm
On the basis of the BLER expression given in (19), we prove that a blocklength can be reduced by grouping semantic bits with the same protection level. This result is formally stated in the following theorem:
@@
Additionally, we formally characterize the condition under which grouping semantic bits results in a longer blocklength. This finding is stated in the following theorem:
@@
Proposition 2 establishes the advantage of grouping semantic bits whose target protection levels are similar. Conversely, Theorem 1 emphasizes the necessity of splitting groups when semantic bits have significantly different protection levels. The left-hand side of (21) quantifies the BLER difference between two groups. Specifically, as $\textsf{BLER}(n_{2},k_{2})$ increases, the ratio $Q^{-1}(\textsf{BLER}(n_{1},k_{1}))/Q^{-1}(\textsf{BLER}(n_{2},k_{2}))$ also increases, making separation of the groups preferable. This intuitive result provides a quantitative criterion for deciding when to merge or separate groups. Motivated by this observation, we propose a grouping algorithm that minimizes the overall blocklength while ensuring that the target bit-flip probabilities are satisfied.
Proposition 2 indicates that grouping bits with comparable error protection requirements reduces blocklength, but it does not explicitly define what constitutes comparable. To address this, we leverage a repetition-based UEP criterion: each semantic bit $i$ is assigned a repetition number $R_{i}$ such that its bit-flip probability remains below the target $\mu_{i}$. Bits sharing the same repetition number inherently possess equivalent protection levels, as they receive identical error protection. We thus group bits according to identical $R_{i}$ values, exploiting coding gain without violating repetition-based UEP constraints. Specifically, the first group includes bits whose repetition numbers are $R_{\max}$, the second group includes bits with repetition number $R_{\max}-2$, and so forth, until the final group contains bits with the minimum repetition number (i.e., $R_{i}=3$). It is worth noting that block channel coding is not applied to semantic bits with a repetition number of one (i.e., $R_{i}=1$), as these bits already meet their target error protection requirements without the need for additional coding.
After this initial grouping, the total number of the non-singleton bit groups is $(\max_{i}R_{i}-1)/2$. To further reduce the overall blocklength, we utilize the criterion derived from Theorem 1. Although Theorem 1 originally provides an equivalent condition favoring separation, it implicitly produces the following condition for merging groups:
@@
However, since the exact blocklengths $n_{1}$ and $n_{2}$ required to achieve the target bit-flip probabilities for groups of sizes $k_{1}$ and $k_{2}$ respectively are unknown, we approximate the BLER of any group $\mathcal{A}$ as follows:
@@
which estimates the probability that at least one bit within a group of size $|\mathcal{A}|$ is erroneous [28]. Using this approximation, we further obtain
@@
allowing us to apply the merging condition effectively. The proposed grouping algorithm is detailed in Algorithm 3.
Algorithm 3 Proposed bit grouping algorithm
@@
24:   $\mathcal{A}\leftarrow\mathcal{J}_{i+1}$
@@
Algorithm 4 Algorithm for fitting group sizes
@@
6:   if $|\mathcal{K}_{g}|<k^{\star}$ and $k_{\rm acc}+k^{\star}\leq K_{\rm group}$ then
@@
18:   else if $|\mathcal{K}_{g}|<k^{\star}$ and $k_{\rm acc}+k^{\star}>K_{\rm group}$ then
@@
IV-C Channel Coding for Block-Level UEP
@@
After determining the $G$ groups $\{\mathcal{K}_{g}\}_{g=1}^{G}$, the semantic bits within each group are jointly encoded by a channel encoder designed to support the minimum target bit-flip probability among the bits in that group. To utilize the capacity-based analysis in Proposition 2 and Theorem 1, which serve as theoretical guidelines under the assumption of capacity-achieving codes, we employ modern channel coding schemes to minimize the gap between theory and practice. Specifically, we employ polar codes and LDPC codes. Polar codes are known to achieve capacity at infinite blocklength and outperforming LDPC in shorter lengths [29], whereas LDPC codes typically exhibit better performance at moderate to high code rates and longer blocklength [30]. Therefore, we apply polar codes and LDPC codes for short and large blocklengths, respectively. Further details on adopted channel coding schemes are provided in Sec. V.
@@
To ensure compatibility with practical polar/LDPC codes and reduce implementation complexity, we constrain the possible group sizes to a predefined set $\overline{\mathcal{K}}$. Note that larger group sizes tend to incur higher decoding complexity due to increased dimensions of the parity-check matrix and the iteration overhead in belief propagation decoding. Furthermore, to maintain tractable memory complexity, we restrict the possible code rates to the set $\mathcal{R}$, e.g. $\mathcal{R}=\{1/2,1/3,1/4\}$, motivated by modulation and coding scheme (MCS) tables commonly found in modern wireless standards [31]. Consequently, we modify the group sizes produced by Algorithm 3 to align with the set $\overline{\mathcal{K}}$. This adjustment process, detailed in Algorithm 4, merges or splits adjacent groups as necessary, ensuring each resulting group $\mathcal{K}_{i}^{\prime}$ closely matches an element of $\overline{\mathcal{K}}$. Specifically, if a group size is smaller than the required code length, subsequent groups are merged to reach the necessary size. Conversely, if a group is excessively large, surplus bits are moved to the following group. This ensures compatibility with practical polar and LDPC codebooks.
@@
To summarize the complete block-level UEP procedure, we present a unified algorithm incorporating the previously discussed grouping and rate-selection steps. As detailed in Algorithm 5, the procedure begins by executing the grouping algorithm (Algorithm 3), followed by adjustments to the group sizes to conform to the allowed group sizes (Algorithm 4). Subsequently, for each group, the most suitable code rate is selected from the predefined set $\mathcal{R}$ by referencing a lookup table $T(k,r)$ that maps group sizes and code rates to achievable bit-flip probabilities. The selected code rate ensures that the bit-flip probability of all bits within the group does not exceed the smallest target bit-flip probability among its members. This integrated approach enables the block-level UEP framework to achieve varying protection levels systematically, while remaining fully compatible with practical codebooks.
@@ V Simulation Results
TABLE I: Semantic encoder and decoder architectures for the MNIST dataset.
Index Layer Name NN Units Activation
Encoder
1 CONV1 32, 3 $\times$ 3 PReLU
2 CONV2 64, 3 $\times$ 3 PReLU
3 CONV3 64, 5 $\times$ 5 PReLU
4 CONV4 8, 5 $\times$ 5 None
5 Flatten - None
Decoder
6 Unflatten - None
7 DECONV1 64, 5 $\times$ 5 PReLU
8 DECONV2 64, 5 $\times$ 5 PReLU
9 DECONV3 32, 3 $\times$ 3 PReLU
10 DECONV4 1, 4 $\times$ 4 None
TABLE II: Semantic encoder and decoder architectures for the CIFAR-10 and CIFAR-100 datasets.
Index Layer Name NN Units Activation
Encoder
1 CONV1 64, 5 $\times$ 5 PReLU
2 CONV2 128, 5 $\times$ 5 PReLU
3 CONV3 128, 5 $\times$ 5 PReLU
4 CONV4 128, 5 $\times$ 5 PReLU
5 CONV5 24, 5 $\times$ 5 None
6 Flatten - None
Decoder
7 Unflatten - None
8 DECONV1 128, 5 $\times$ 5 PReLU
9 DECONV2 128, 5 $\times$ 5 PReLU
10 DECONV3 128, 5 $\times$ 5 PReLU
11 DECONV4 64, 5 $\times$ 5 PReLU
12 DECONV5 3, 5 $\times$ 5 None
The encoder–decoder pair and the bit-flip probabilities $\{\mu_{i}\}$ are jointly trained as defined in (8). We employ convolutional neural network (CNN)-based autoencoder architectures, following the design in [6], which are detailed in Table I and Table II. The feature vector is quantized using 8 bits. The symbol transmission power is set to $P_{\rm trans}=0$ dBW, and the SNR is fixed at 0 dB under an AWGN channel.
@@ 12 DECONV5 3, 5 $\times$ 5 None
  • Block-UEP: This is the proposed block-level UEP framework. Using the learned $\{\mu_{i}\}$, we apply Algorithm 5 to group bits with similar protection requirements. The set of allowable group sizes is $\mathcal{K}=\{128,\,256,\,512,\,1024\}$. Each group is encoded using either a polar code (for group sizes of 128 and 256) or an LDPC code (for group sizes of 512 and 1024), where the code rate is selected from a lookup table with available code rates $\mathcal{R}=\{3/4,\,2/3,\,15/24,\,14/24,\,13/24,\,1/2,\,1/3\}$. Both polar and LDPC codes are implemented according to the 5G standard: polar codes utilize CRC-aided list decoding with a list size of 8, while LDPC codes are constructed using the standard’s two base graphs and a lifting procedure [31].
@@
  • LDPC: This is an equal error protection scheme using an LDPC code. Bit grouping is performed identically to the Block-UEP scheme. However, all groups are encoded using LDPC codes with a fixed rate, which is selected to ensure that the total blocklength matches that of the Block-UEP scheme.
  • Polar: This is an equal error protection scheme using a polar code. Since a large group size is prohibitive due to polar code’s implementation complexity, all semantic bit groups are uniformly set to size 256. All groups use the same code rate selected to ensure that the total blocklength matches that of the Block-UEP scheme.
@@
In Figs. 5 and 6, we compare the PSNR and corresponding total blocklength for various channel coding frameworks, using the CIFAR-10 dataset with $\lambda=10^{-4}$ and $\lambda=10^{-6}$, respectively. The results demonstrate that the proposed bit-level UEP framework (Bit-UEP) consistently achieves higher PSNR compared to fixed-rate repetition coding (Repetition) under both training settings. Specifically, the fixed-rate scheme with $R_{\rm fix}=8$ consumes slightly more blocklength than Bit-UEP, yet suffers from a significantly lower PSNR. On the other hand, the fixed-rate scheme with $R_{\rm fix}=12$ achieves comparable PSNR to Bit-UEP but at the cost of significantly larger blocklength usage. Moreover, coding gain is evident when comparing Bit-UEP and Block-UEP. Block-UEP achieves higher PSNR with even less blocklength usage compared to Bit-UEP. Additionally, Block-UEP outperforms traditional equal rate schemes (LDPC and Polar) in terms of PSNR while maintaining the same blocklength across both datasets. This result demonstrates the importance and effectiveness of dynamic grouping of the semantic bits according to their protection requirements.
Similarly, in Figs. 7 and 8, we compare the PSNR and total blocklegnth across various channel coding frameworks using the CIFAR-100 dataset, with neural networks trained under $\lambda=10^{-4}$ and $\lambda=10^{-6}$, respectively. The results again confirm the superiority of unequal error protection over traditional equal protection schemes. From the consistent observations on both CIFAR-10 and CIFAR-100, we conclude that the proposed UEP frameworks provide clear advantages for semantic communication in terms of both reconstruction quality and transmission efficiency. Notably, the performance of Block-UEP approaches the genie-aided upper bound in both datasets, indicating its effectiveness in achieving near-optimal reconstruction quality with an improved transmission efficiency. In Fig. 9, we visualize this effect with reconstructed CIFAR-100 images. The figure demonstrates that Block-UEP offers the highest image quality, while the equal error protection schemes (LDPC and Polar) suffer from image quality degradation.
@@
TABLE III: PSNR of the proposed block-level UEP framework with different orderings of semantic bits.
@@
In Table III, we compare the PSNR of the proposed block-level UEP framework with different orderings of semantic bits. In the proposed ordering, the semantic bits are sorted in ascending order of their target bit-flip probabilities, as described in Sec. III and Sec. IV, and Algorithm 5 is applied to this sorted sequence. For comparison, we consider two alternative orderings: random and reverse. As shown in Table III, the proposed ordering achieves the highest PSNR, while the reverse ordering yields the worst performance. This result confirms that bits with lower target bit-flip probabilities are more critical for image reconstruction. In Fig. 10, we also visualize this effect with reconstructed CIFAR-10 images. The figure demonstrates that reversing the bit order leads to severe degradation in image quality.
In Figs. 11 and 12, we compare the PSNR and total blocklength of the proposed block-level UEP framework with different values of $\lambda$ using the MNIST and CIFAR-10 datasets, respectively. The regularization parameter $\lambda$ penalizes the assignment of excessively low bit-flip probabilities to task-irrelevant bits in the loss function (8). As shown in the figures, a larger $\lambda$ tends to reduce PSNR, as it limits the model’s capacity to minimize reconstruction error. However, it also significantly decreases total blocklength in both datasets. These results suggest that selecting an appropriate value of $\lambda$ allows for a trade-off between reconstruction quality and transmission efficiency, depending on the desired task performance level.

### v2

@@ IV-B Theoretical Foundation for Block Partitioning
On the basis of the BLER expression given in (20), we prove that a blocklength can be reduced by partitioning semantic bits into blocks of the same protection level. This result is formally stated in the following theorem:
@@
Additionally, we formally characterize the condition under which block partitioning results in a longer blocklength. This finding is stated in the following theorem:
@@
IV-C Proposed Block Partitioning Algorithm
Proposition 2 establishes the advantage of partitioning semantic bits with similar target protection levels into common blocks. Conversely, Theorem 1 emphasizes the necessity of splitting blocks when semantic bits have significantly different protection levels. The left-hand side of (22) quantifies the BLER difference between two blocks. Specifically, as $\textsf{BLER}(n_{2},k_{2})$ increases, the ratio $Q^{-1}(\textsf{BLER}(n_{1},k_{1}))/Q^{-1}(\textsf{BLER}(n_{2},k_{2}))$ also increases, making separation of the blocks preferable. This intuitive result provides a quantitative criterion for deciding when to merge or separate blocks. Motivated by this observation, we propose a block partitioning algorithm that minimizes the overall blocklength while ensuring that the target bit-flip probabilities are satisfied.
Proposition 2 indicates that partitioning comparable error protection requirements into common blocks reduces blocklength, but it does not explicitly define what constitutes comparable. To address this, we leverage a repetition-based UEP criterion: each semantic bit $i$ is assigned a repetition number $R_{i}$ such that its bit-flip probability remains below the target $\mu_{i}$. Bits sharing the same repetition number inherently possess equivalent protection levels, as they receive identical error protection. We thus partition bits according to identical $R_{i}$ values, exploiting coding gain without violating repetition-based UEP constraints. Specifically, the first block includes bits whose repetition numbers are $R_{\max}$, the second block includes bits with repetition number $R_{\max}-2$, and so forth, until the final block contains bits with the minimum repetition number (i.e., $R_{i}=3$). It is worth noting that block channel coding is not applied to semantic bits with a repetition number of one (i.e., $R_{i}=1$), as these bits already meet their target error protection requirements without the need for additional coding.
After this initial partitioning, the total number of the non-singleton bit blocks is $(\max_{i}R_{i}-1)/2$. To further reduce the overall blocklength, we utilize the criterion derived from Theorem 1. Although Theorem 1 originally provides an equivalent condition favoring separation, it implicitly produces the following condition for merging blocks:
@@ IV-C Proposed Block Partitioning Algorithm
However, since the exact blocklengths $n_{1}$ and $n_{2}$ required to achieve the target bit-flip probabilities for blocks of sizes $k_{1}$ and $k_{2}$ respectively are unknown, we approximate the BLER of any block $\mathcal{A}$ as follows:
@@
which estimates the probability that at least one bit within a block of size $|\mathcal{A}|$ is erroneous [21]. Using this approximation, we further obtain
@@
allowing us to apply the merging condition effectively. The proposed block partitioning algorithm is detailed in Algorithm 3.
Algorithm 3 Proposed block paritioning algorithm
@@
24:   $\mathcal{A}\leftarrow\mathcal{J}_{p+1}$
@@
Algorithm 4 Algorithm for fitting block sizes
@@
6:   if $|\mathcal{K}_{g}|<k^{\star}$ and $k_{\rm acc}+k^{\star}\leq K_{\rm block}$ then
@@
18:   else if $|\mathcal{K}_{g}|<k^{\star}$ and $k_{\rm acc}+k^{\star}>K_{\rm block}$ then
@@
IV-D A Systematic Approach to Incorporating Modern Channel Codes
@@
After determining the $G$ blocks $\{\mathcal{K}_{g}\}_{g=1}^{G}$, the semantic bits within each block can be jointly encoded by a channel encoder. To utilize the capacity-based analysis in Proposition 2 and Theorem 1, which serve as theoretical guidelines under the assumption of capacity-achieving codes, we employ modern channel coding schemes to minimize the gap between theory and practice. Specifically, we employ polar codes and LDPC codes. Polar codes are known to achieve capacity at infinite blocklength and outperforming LDPC in shorter lengths [31], whereas LDPC codes typically exhibit better performance at moderate to high code rates and longer blocklength [27]. Therefore, we apply polar codes and LDPC codes for short and large blocklengths, respectively. Further details on adopted channel coding schemes are provided in Sec. V.
@@
To ensure compatibility with practical polar/LDPC codes and reduce implementation complexity, we constrain the possible block sizes to a predefined set $\overline{\mathcal{K}}$. Note that larger block sizes tend to incur higher decoding complexity due to increased dimensions of the parity-check matrix and the iteration overhead in belief propagation decoding. Furthermore, to maintain tractable memory complexity, we restrict the possible code rates to the set $\mathcal{R}$, e.g. $\mathcal{R}=\{1/2,1/3,1/4\}$, motivated by modulation and coding scheme (MCS) tables commonly found in modern wireless standards [1]. Consequently, we modify the block sizes produced by Algorithm 3 to align with the set $\overline{\mathcal{K}}$. This adjustment process, detailed in Algorithm 4, merges or splits adjacent blocks as necessary, ensuring each resulting block $\mathcal{K}_{i}^{\prime}$ closely matches an element of $\overline{\mathcal{K}}$. Specifically, if a block size is smaller than the required code length, subsequent blocks are merged to reach the necessary size. Conversely, if a block is excessively large, surplus bits are moved to the following block. This ensures compatibility with practical polar and LDPC codebooks.
@@
To summarize the complete block-level UEP procedure, we present a unified algorithm incorporating the previously discussed block partitioning and rate-selection steps. As detailed in Algorithm 5, the procedure begins by executing the block partitioning algorithm (Algorithm 3), followed by adjustments to the block sizes to conform to the allowed block sizes (Algorithm 4). Subsequently, for each block, the most suitable code rate is selected from the predefined set $\mathcal{R}$ by referencing a lookup table $T(k,r)$ that maps block sizes and code rates to achievable bit-flip probabilities. The selected code rate ensures that the bit-flip probability of all bits within the block does not exceed the smallest target bit-flip probability among its members. This integrated approach enables the block-level UEP framework to achieve varying protection levels systematically, while remaining fully compatible with practical codebooks.
@@ V Simulation Results
TABLE I: The semantic encoder and decoder architectures for MNIST, CIFAR-10, and CIFAR-100 datasets.
Layers
MNIST Encoder C(32,3,1,2), PReLU, C(64,3,1,2), PReLU,
C(64,5,2,1), PReLU, C(8,5,2,1)
Decoder CT(64,5,2,1,0), PReLU, CT(64,5,2,1,0), PReLU,
CT(32,3,1,2,0), PReLU, CT(1,4,1,2,0)
Encoder C(64,5,2,2), PReLU, C(128,5,2,2), PReLU,
CIFAR-10 C(128,5,1,2), PReLU, C(128,5,1,2), PReLU,C(24,5,1,2)
/ CIFAR-100 Decoder CT(128,5,1,2,0), PReLU, CT(128,5,1,2,0), PReLU,
CT(128,5,1,2,0), PReLU, CT(64,5,2,2,1), PReLU,CT(3,5,2,2,1)
The encoder–decoder pair and the bit-flip probabilities $\{\mu_{i}\}$ are jointly trained as defined in (9). We employ convolutional neural network (CNN)-based autoencoder architectures, following the design in [4], which are detailed in Table I. In this table, ${\rm C}(c,k,s,p)$ represents a 2D convolutional layer with $c$ output channels, a $k\times k$ kernel, a stride of $s$, and padding of $p$. Similarly, ${\rm CT}(c,k,s,p,p_{o})$ denotes a transposed 2D convolutional layer with output padding of $p_{0}$. The feature vector is quantized using 8 bits. The symbol transmission power is set to $P_{\rm trans}=0$ dBW, and the received SNR defined as $|h|^{2}/2\sigma^{2}$ is fixed at 0 dB under the Rayleigh block fading channel.
@@
  • Block-UEP: This is the proposed block-level UEP framework. Using the learned $\{\mu_{i}\}$, we apply Algorithm 5 to perform block partitioning. The set of allowable block sizes is $\mathcal{K}=\{128,\,256,\,512,\,1024\}$. Each block is encoded using either a polar code (for block sizes of 128 and 256) or an LDPC code (for block sizes of 512 and 1024), where the code rate is selected from a lookup table with available code rates $\mathcal{R}=\{3/4,\,2/3,\,15/24,\,14/24,\,13/24,\,1/2,\,1/3\}$. Both polar and LDPC codes are implemented according to the 5G standard: polar codes utilize CRC-aided list decoding with a list size of 8, while LDPC codes are constructed using the standard’s two base graphs and a lifting procedure [1].
@@
  • LDPC: This is an equal error protection scheme using an LDPC code. Block partitioning is performed identically to the Block-UEP scheme. However, all blocks are encoded using LDPC codes with a fixed rate, which is selected to ensure that the total blocklength matches that of the Block-UEP scheme.
  • Polar: This is an equal error protection scheme using a polar code. Since a large block size is prohibitive due to polar code’s implementation complexity, all semantic bit blocks are uniformly set to size 256. All blocks use the same code rate selected to ensure that the total blocklength matches that of the Block-UEP scheme.
@@
In Figs. 4 and 5, we compare the PSNR and corresponding total blocklength for various channel coding frameworks, using the CIFAR-10 dataset with $\lambda=10^{-4}$ and $\lambda=10^{-6}$, respectively. The results demonstrate that the proposed bit-level UEP framework (Bit-UEP) consistently achieves higher PSNR compared to fixed-rate repetition coding (Repetition) under both training settings. Specifically, the fixed-rate scheme with $R_{\rm fix}=8$ consumes slightly more blocklength than Bit-UEP, yet suffers from a significantly lower PSNR. On the other hand, the fixed-rate scheme with $R_{\rm fix}=12$ achieves comparable PSNR to Bit-UEP but at the cost of significantly larger blocklength usage. Moreover, coding gain is evident when comparing Bit-UEP and Block-UEP. Block-UEP achieves higher PSNR with even less blocklength usage compared to Bit-UEP. Additionally, Block-UEP outperforms traditional equal rate schemes (LDPC and Polar) in terms of PSNR while maintaining the same blocklength across both datasets. This result demonstrates the importance and effectiveness of dynamic partitioning of the semantic bits according to their protection requirements.
Similarly, in Figs. 6 and 7, we compare the PSNR and total blocklength across various channel coding frameworks using the CIFAR-100 dataset, with $\lambda=10^{-4}$ and $\lambda=10^{-6}$, respectively. The results again confirm the superiority of unequal error protection over traditional equal protection schemes. From the consistent observations on both CIFAR-10 and CIFAR-100, we conclude that the proposed UEP frameworks provide clear advantages for semantic communication in terms of both reconstruction quality and transmission efficiency. Notably, the performance of Block-UEP approaches the genie-aided upper bound in both datasets, indicating its effectiveness in achieving near-optimal reconstruction quality with an improved transmission efficiency. In Fig. 8, we visualize this effect with reconstructed CIFAR-100 images. The figure demonstrates that Block-UEP offers the highest image quality, while the equal error protection schemes (LDPC and Polar) suffer from image quality degradation.
@@
In Table II, we compare the PSNR of the proposed block-level UEP framework with different orderings of semantic bits. In the proposed ordering, the semantic bits are sorted in ascending order of their target bit-flip probabilities, as described in Sec. III and Sec. IV, and Algorithm 5 is applied to this sorted sequence. For comparison, we consider two alternative orderings: random and reverse. As shown in Table II, the proposed ordering achieves the highest PSNR, while the reverse ordering yields the worst performance. This result confirms that bits with lower target bit-flip probabilities are more critical for image reconstruction. In Fig. 9, we also visualize this effect with reconstructed CIFAR-10 images. The figure demonstrates that reversing the bit order leads to severe degradation in image quality.
TABLE II: PSNR of the proposed block-level UEP framework with different orderings of semantic bits.
@@
In Figs. 10 and 11, we compare the PSNR and total blocklength of the proposed block-level UEP framework with different values of $\lambda$ using the MNIST and CIFAR-10 datasets, respectively. The regularization parameter $\lambda$ penalizes the assignment of excessively low bit-flip probabilities to task-irrelevant bits in the loss function (9). As shown in the figures, a larger $\lambda$ tends to reduce PSNR, as it limits the model’s capacity to minimize reconstruction error. However, it also significantly decreases total blocklength in both datasets. These results suggest that selecting an appropriate value of $\lambda$ allows for a trade-off between reconstruction quality and transmission efficiency, depending on the desired task performance level.
VI Conclusion
In this paper, we explored the UEP problem in digital semantic communication by leveraging learned bit-flip probabilities as target protection levels. We proposed two coding frameworks: a bit-level UEP strategy based on repetition coding and a block-level UEP approach using polar codes and LDPC codes. The bit-level framework minimizes total blocklength while ensuring per-bit reliability, whereas the block-level framework enhances coding efficiency by partitioning bits with similar protection requirements into common blocks. For the block-level framework, we introduced a block partitioning algorithm guided by finite blocklength capacity analysis. Simulation results demonstrated that both frameworks significantly improve task performance and transmission efficiency compared to conventional methods. These results highlight the importance of bit-level error protection in semantic communication and lay the groundwork for future research on resource-efficient and reliability-aware semantic transmission strategies.